OpenStudy (anonymous):

Find the Lim x->0 ((1/x+2)-(1/2))/(x)

OpenStudy (anonymous):

\[limx-0 \frac{\frac{1}{x+2}-\frac{1}{2}}{x}\]

OpenStudy (anonymous):

is that the question?

OpenStudy (anonymous):

Yes. Sorry I dont know how to make it look like that yet.

OpenStudy (anonymous):

no probs, start by combininb the top fractions, so get common denominators \[\lim x-0 \frac{\frac{2-1(x+2)}{2(x+2)}}{x}\]

OpenStudy (anonymous):

okay I have that

OpenStudy (anonymous):

Wait..I dont get it. the thing you just posted confused me.

OpenStudy (zarkon):

\[\frac{\frac{2-1(x+2)}{2(x+2)}}{x}=\frac{1}{x}\frac{2-1(x+2)}{2(x+2)}\]

OpenStudy (anonymous):

yes i had to delete, inversed x accidentally, cancel your -x on top with the bottom \[limx-0 \frac{-1}{2(x+2}\]

OpenStudy (zarkon):

\[\LaTeX\] \[\lim_{x\to a}\] \lim_{x\to a}

OpenStudy (anonymous):

how did you get -1 on the top?

OpenStudy (anonymous):

thanks zarkon

OpenStudy (anonymous):

\[\frac{1}{x} * \frac{-x}{2(x+2)}\]

OpenStudy (anonymous):

the x's cancel

OpenStudy (anonymous):

okay but where did the (x+2) go? 2-1(x+2)

OpenStudy (zarkon):

\[\frac{1}{x} \cdot \frac{-x}{2(x+2)}=\frac{1}{x} \cdot \frac{(-1)x}{2(x+2)}\] \[=\frac{1}{\cancel{x}} \cdot \frac{(-1)\cancel{x}}{2(x+2)}=\frac{-1}{2(x+2)}\]

OpenStudy (zarkon):

\[2-1(x+2)=2-x-2=-x\]

OpenStudy (anonymous):

Oooh Okay. I got it now. I accidentaly distributed before. The answer is -1/4 when you plug in the zero right?

OpenStudy (zarkon):

yes

OpenStudy (anonymous):

This looks somehow like the derived formula of the first principle to me. :P Ok back to solving. \[\lim_{x \rightarrow 0} ((1/(x+2))-1/2)/x\]\[=\lim_{x \rightarrow 0} (2-(x+2))/(2(x+2))/x\]\[=\lim_{x \rightarrow 0} (-x)/(2(x+2))/x\]\[=\lim_{x \rightarrow 0} -1/(2x+4)\]\[=-1/4\]

OpenStudy (anonymous):

Zarkon: two latex questions if you don't mind: 1. How do you do that canceling out (the diagonal lines on top of the X's) ? 2. How did write that fancy "Latex" in your second post ? Meanwhile, looking for good reference on Latex (I see many on google). Thanx!

OpenStudy (zarkon):

\[\cancel{Zarkon}\] \cancel{Zarkon} \[\LaTeX\] \LaTeX

OpenStudy (anonymous):

Thank you Zarkon - although there was no need to cancel yourself for the demonstration :)